# envname-P envname#1

### v2

@@ Abstract
While large language models (LLMs) with reasoning capabilities are progressing rapidly on high-school math competitions and coding, can they reason effectively through complex, open-ended challenges found in frontier physics research? And crucially, what kinds of reasoning tasks do physicists actually want LLMs to assist with? To address these questions, we present the CritPt (Complex Research using Integrated Thinking - Physics Test, pronounced “critical point”), the first benchmark designed to test LLMs on unpublished, research-level reasoning tasks that broadly covers modern physics research areas, including condensed matter, quantum physics, atomic, molecular & optical physics, astrophysics, high energy physics, mathematical physics, statistical physics, nuclear physics, nonlinear dynamics, fluid dynamics and biophysics. CritPt consists of 71 composite research challenges designed to simulate full-scale research projects at the entry level, which are also decomposed to 190 simpler checkpoint tasks for more fine-grained insights. All problems are newly created by over 50 active physics researchers based on their own research. Every problem is hand-curated to admit a guess-resistant and machine-verifiable answer and is evaluated by an automated grading pipeline heavily customized for advanced physics-specific output formats. We find that while current state-of-the-art LLMs show early promise on isolated checkpoints, they remain far from being able to reliably solve full research-scale challenges: the best average accuracy among base models is only 4.0% , achieved by GPT-5 (high), moderately rising to around 10% when equipped with coding tools. Through the realistic yet standardized evaluation offered by CritPt, we highlight a large disconnect between current model capabilities and realistic physics research demands, offering a foundation to guide the development of scientifically grounded AI tools.
@@ 1 Introduction
In Sec. 4, we show that current state-of-the-art LLMs are making early progress on isolated checkpoints, but remain far from being able to reliably solve full research-scale challenges. Even the best-performing base model on CritPt, GPT-5 (high), reaches only 4.0% average accuracy on challenges, with most other models scoring near zero. When equipped with tools (code interpreter and web search), GPT-5 (high) improves modestly to 11.7% accuracy. More stringent evaluation metrics reveal that current LLMs still lack reliability when tackling research-level problems, underscoring the gap between today’s models and the demands of realistic physics research workflow.
@@ 2 Design choices of CritPt
We begin by describing the data sources and coverage of CritPt in Sec. 2.1, followed by the technical problem criteria in Sec. $2.$ The data creation and review process are outlined in Sec. 2.3, and Sec. 2.4 presents the structure of a CritPt challenge with an illustrative example.
@@ 2.1 Source and coverage: hand-curated research challenges from the physics community
Research Discipline Challenges % of Total Checkpoints % of Total
Condensed Matter Physics 25 35.2% 69 36.3%
Atomic, Molecular & Optical 15 21.1% 42 22.1%
Quantum Information, Science & Technology 15 21.1% 39 20.5%
Gravitation, Cosmology & Astrophysics 11 15.5% 30 15.8%
@@
Mathematical Physics 9 12.7% 20 10.5%
@@
Table 1: The physics research disciplines covered by CritPt’s challenges and checkpoints.
@@ 4 Results
In this section, we evaluate 10 state-of-art models (configuration in Table 2) to directly answer the three lines of inquiry in the introduction. Evaluation is performed independently on two levels: the full challenges (Sec. 4.1) and finer-grained checkpoints (Sec. 4.2). In these two sections, we report average accuracy over five runs as our primary metric to reduce high stochasticity in model behavior. Next in Sec. 4.3, we adopt a stricter metric, consistently solved rate, to probe reliability of model performance. Table 3 is a brief summary of these aggregated accuracy statistics. Detailed analysis of model responses beyond aggregated statistics is discussed in Sec. $4.$
@@
GPT-5 (high, code & web) 11.7 20.8 25.6 8.6 15.0 18.7
GPT-5 (high, code) 09.4 18.7 24.6 5.7 15.0 18.7
GPT-5 (high) 04.0 14.4 20.5 2.9 09.6 16.0
Gemini-2.5 Pro 01.7 07.4 09.6 0.0 03.7 04.8
o3 (high) 01.1 07.3 11.0 0.0 03.2 06.4
Gemini-2.5 Flash 01.1 02.8 04.8 0.0 01.7 02.7
DeepSeek R1 00.9 04.8 06.5 0.0 01.7 03.2
o4-mini (high) 00.6 05.8 07.9 0.0 02.7 03.2
Claude Opus 4 00.3 02.8 04.2 0.0 00.5 02.1
GPT-5 (minimal) 00.0 02.8 04.5 0.0 01.6 02.7
Llama-4 Maverick 00.0 02.6 02.9 0.0 01.1 01.1
GPT-4o 00.0 01.8 01.6 0.0 00.0 00.0
@@ 4.1 Challenge-level evaluation: can LLMs solve unseen research problems?
We first assess model performance on CritPt challenges without intermediate supervision, testing their end-to-end complex reasoning in unseen full-scale research problems. As shown in Fig. 3, all models score low in terms of average accuracy of five independent runs across 70 challenges in CritPt test set. Among base models (no external tools), GPT-5 (high), achieves only 4.0% as the best result, with all other models under 2%.
@@
Tool use can offer modest but meaningful gains. With access to the code interpreter, GPT-5 (high, code) improves to 9.4%, a notable relative improvement from the base model, consistent with the importance of numerical tools in modern physics research. Adding web search brings only a marginal increase to 11.7%, confirming that CritPt’s search-proof design effectively resists short-cutting via retrieval and emphasizes genuine reasoning.
@@ 4.2 Checkpoint-level evaluation: smaller tasks that LLMs can assist today?
As shown in Fig. 5, current LLMs show early promise in assisting with checkpoints, more localized or well-scoped tasks compared to challenges. GPT-5 (high) again outperforms all other models. In the self-carryover without expert answer (solid bar), GPT-5 (high) reaches 14.4%. This improves to 18.7% with code interpreter and reaches 20.8% when also equipped with web search. The next best performers are Gemini 2.5 Pro (7.4%) and o3 (high) (7.3%), followed by o4-mini (high) (5.8%) and DeepSeek R1 (4.8%). The remaining models fall below 3%, where small absolute differences may reflect statistical noise rather than meaningful performance gaps, so the ordering should be interpreted with caution.
Almost all models benefit from expert answer injection to earlier checkpoints. The largest gains are observed for GPT-5 (high) family and o3 (high), which improve by more than 3.7% in the oracle-carryover setting (hatched bar), indicating the ability to leverage correct intermediate results to improve downstream reasoning.
@@ 4.3 Reliability metric: can we trust LLM outputs?
At the full challenge level (Fig. 6a), only GPT-5 (high) is able to solve any problems consistently [67]. The base model scores only 2.9% (2 out of 70 challenges). With tool use, this improves to 5.7% (code) and 8.6% (code & web). All other reasoning-oriented models drop to zero under this stricter measure.
At the checkpoint level (Fig. 6b), leading models are able to consistently solve a very limited number. Interestingly, GPT-5 (high, code) and GPT-5 (high, code & web) achieve identical scores here, further supporting the search-proof design of CritPt even at the checkpoint level.
@@ 4.4 Detailed analysis of full model responses
Beyond aggregated accuracy metrics, we further analyze model behavior at the level of individual challenges, and walk through the full model responses (including reasoning traces when available) with physics experts, which helps surface more qualitative insights not captured by final answer correctness alone.
To support an efficient review process with experts and overcome the limitations of traditional LLM infrastructure, we develop an interactive visualization platform. This tool allows experts to quickly browse model responses, systematically compare performance across tasks and model families, and identify error patterns or interesting behaviors. Table 4 illustrates the visualization of the example challenge, where we can immediately observe an unexpected behavior: adding web search tool on top of coding for GPT-5 (high) actually degrades model performance for this particular challenge. The full interactive demo including all model responses is available at critpt.com/example.
@@ A.1 List of CritPt challenges
  36. 36.
    Dark matter with a dark photon mediator
@@
    Scattering rate of the HK model
@@
    Axion in neutron stars
@@
    Scalar DM in cosmic explorer
@@
    Vector DM in modified LIGO
@@
    Magnetic space group identification
@@
    Optical Stark shift and Bloch-Siegert shift
@@
    Superresolution
@@
    Quantum search time
@@
    Quantum games in multi-slit interference
@@
    Noise robustness in Kochen–Specker and Spekkens contextuality
@@
    Constructing fermionic matrix operators
@@
    Quantum capacity for quantum channels
@@
    Quantum $f$-divergence
@@ A.4 API usage statistics
GPT-5 (high, code & web) 328157 1.25 288301 0.125 10413 10.0 2297 10.0 340867 0.573
GPT-5 (high, code) 17960 1.25 16784 0.125 11120 10.0 1999 10.0 31078 0.156
GPT-5 (high) 856 1.25 63 0.125 24834 10.0 2048 10.0 27738 0.270
o3 (high) 856 2.0 91 0.5 17858 8.0 1435 8.0 20150 0.156
o4-mini (high) 856 1.1 122 0.275 16501 4.4 1534 4.4 18892 0.080
Gemini 2.5 Pro 996 1.25 0 0.31 20984 10.0 3226 10.0 25205 0.243
Gemini 2.5 Flash 996 0.3 0 0.075 22523 2.5 3799 2.5 27318 0.066
DeepSeek R1 814 0.56 339 0.07 19572 1.68 1241 1.68 21627 0.035
Claude Opus 4 996 15.0 0 1.5 4588 75.0 8728 75.0 14312 1.014
GPT-5 (minimal) 856 1.25 20 0.125 0 10.0 2294 10.0 3150 0.024
GPT-4o 857 2.5 212 1.25 0 10.0 952 10.0 1809 0.012
Llama-4 Maverick 810 0.27 0 0.0 0 0.85 1383 0.85 2194 0.001

### v3

@@ Abstract
While large language models (LLMs) with reasoning capabilities are progressing rapidly on high-school math competitions and coding, can they reason effectively through complex, open-ended challenges found in frontier physics research? And crucially, what kinds of reasoning tasks do physicists actually want LLMs to assist with? To address these questions, we present the CritPt (Complex Research using Integrated Thinking - Physics Test, pronounced “critical point”), the first benchmark designed to test LLMs on unpublished, research-level reasoning tasks that broadly covers modern physics research areas, including condensed matter, quantum physics, atomic, molecular & optical physics, astrophysics, high energy physics, mathematical physics, statistical physics, nuclear physics, nonlinear dynamics, fluid dynamics and biophysics. CritPt consists of 71 composite research challenges designed to simulate full-scale research projects at the entry level, which are also decomposed to 190 simpler checkpoint tasks for more fine-grained insights. All problems are newly created by over 50 active physics researchers based on their own research. Every problem is hand-curated to admit a guess-resistant and machine-verifiable answer and is evaluated by an automated grading pipeline heavily customized for advanced physics-specific output formats. We find that while current state-of-the-art LLMs show early promise on isolated checkpoints, they remain far from being able to reliably solve full research-scale challenges: the best average accuracy among base models is only 5.7% , achieved by GPT-5 (high), moderately rising to around 10% when equipped with coding tools. Through the realistic yet standardized evaluation offered by CritPt, we highlight a large disconnect between current model capabilities and realistic physics research demands, offering a foundation to guide the development of scientifically grounded AI tools.
@@ 1 Introduction
In Sec. 4, we show that current state-of-the-art LLMs are making early progress on isolated checkpoints, but remain far from being able to reliably solve full research-scale challenges. Even the best-performing base model on CritPt, GPT-5 (high), reaches only 5.7% average accuracy on challenges, with most other models scoring near zero. When equipped with tools (code interpreter and web search), GPT-5 (high) improves modestly to 12.6% accuracy. More stringent evaluation metrics reveal that current LLMs still lack reliability when tackling research-level problems, underscoring the gap between today’s models and the demands of realistic physics research workflow.
@@ 2 Design choices of CritPt
We begin by describing the data sources and coverage of CritPt in Sec. 2.1, followed by the technical problem criteria in Sec. 2.2. The data creation and review process are outlined in Sec. 2.3, and Sec. 2.4 presents the structure of a CritPt challenge with an illustrative example.
@@ 2.1 Source and coverage: hand-curated research challenges from the physics community
Research Area Challenges % of Total Checkpoints % of Total
Condensed Matter Physics 25 35.2% 70 36.8%
Quantum Information, Science & Technology 17 23.9% 43 22.6%
Atomic, Molecular & Optical 14 19.7% 42 22.1%
@@
Mathematical Physics 10 14.1% 22 11.6%
Gravitation, Cosmology & Astrophysics 9 12.7% 25 13.2%
@@
Table 1: The physics research areas covered by CritPt’s challenges and checkpoints.
@@ 4 Results
In this section, we evaluate 10 state-of-art models (configuration in Table 2) to directly answer the three lines of inquiry in the introduction. Evaluation is performed independently on two levels: the full challenges (Sec. 4.1) and finer-grained checkpoints (Sec. 4.2). In these two sections, we report average accuracy over five runs as our primary metric to reduce high stochasticity in model behavior. Next in Sec. 4.3, we adopt a stricter metric, consistently solved rate, to probe reliability of model performance. Table 3 is a brief summary of these aggregated accuracy statistics. Detailed analysis of model responses beyond aggregated statistics is discussed in Sec. 4.4.
@@
GPT-5 (high, code & web) 12.6 21.4 24.5 10.0 14.4 17.6
GPT-5 (high, code) 10.6 20.0 23.9 08.6 15.5 17.6
GPT-5 (high) 05.7 15.3 20.0 04.3 10.7 15.5
Gemini-2.5 Pro 02.0 08.1 09.1 00.0 04.3 03.7
o3 (high) 01.4 07.4 10.6 00.0 03.7 05.9
DeepSeek R1 01.1 05.1 06.6 00.0 02.1 03.2
Gemini-2.5 Flash 01.1 04.1 04.6 00.0 02.1 02.1
o4-mini (high) 00.6 05.6 07.1 00.0 02.7 02.1
Claude Opus 4 00.3 02.6 03.0 00.0 00.5 01.6
GPT-5 (minimal) 00.0 02.6 03.7 00.0 01.6 02.1
Llama-4 Maverick 00.0 02.2 02.1 00.0 01.1 00.5
GPT-4o 00.0 01.7 01.2 00.0 00.0 00.0
@@ 4.1 Challenge-level evaluation: can LLMs solve unseen research problems?
We first assess model performance on CritPt challenges without intermediate supervision, testing their end-to-end complex reasoning in unseen full-scale research problems. As shown in Fig. 3, all models score low in terms of average accuracy of five independent runs across 70 challenges in CritPt test set. Among base models (no external tools), GPT-5 (high), achieves only 5.7% as the best result, with all other models below or equal to 2%.
@@
Tool use can offer modest but meaningful gains. With access to the code interpreter, GPT-5 (high, code) improves to 10.6%, a notable relative improvement from the base model, consistent with the importance of numerical tools in modern physics research. Adding web search brings only a marginal increase to 12.6%, confirming that CritPt’s search-proof design effectively resists short-cutting via retrieval and emphasizes genuine reasoning.
@@ 4.2 Checkpoint-level evaluation: smaller tasks that LLMs can assist today?
As shown in Fig. 5, current LLMs show early promise in assisting with checkpoints, more localized or well-scoped tasks compared to challenges. GPT-5 (high) again outperforms all other models. In the self-carryover without expert answer (solid bar), GPT-5 (high) reaches 15.3%. This improves to 20.0% with code interpreter and reaches 21.4% when also equipped with web search. The next best performers are Gemini 2.5 Pro (8.1%) and o3 (high) (7.4%), followed by o4-mini (high) (5.6%), DeepSeek R1 (5.1%) and Gemini 2.5 Flash (4.1%). The remaining models fall below 3%, where small absolute differences may reflect statistical noise rather than meaningful performance gaps, so the ordering should be interpreted with caution.
Almost all models benefit from expert answer injection to earlier checkpoints. The largest gains are observed for GPT-5 (high) family and o3 (high), which improve by more than 3% in the oracle-carryover setting (hatched bar), indicating the ability to leverage correct intermediate results to improve downstream reasoning.
@@ 4.3 Reliability metric: can we trust LLM outputs?
At the full challenge level (Fig. 6a), only GPT-5 (high) is able to solve any problems consistently [67]. The base model scores only 4.3% (3 out of 70 challenges). With tool use, this improves to 8.6% (code) and 10.0% (code & web). All other reasoning-oriented models drop to zero under this stricter measure.
At the checkpoint level (Fig. 6b), leading models are able to consistently solve a very limited number. Interestingly, GPT-5 (high, code) outperforms GPT-5 (high, code & web) with a small margin, further supporting the search-proof design of CritPt even at the checkpoint level.
@@ 4.4 Detailed analysis of full model responses
Beyond aggregated accuracy metrics, we further analyze model behavior at the level of individual challenges and walk through the full model responses (including reasoning traces when available) with physics experts, which helps surface more qualitative insights not captured by final answer correctness alone.
To support an efficient review process with experts and overcome the limitations of traditional LLM infrastructure, we develop an interactive visualization platform. This tool allows experts to quickly browse model responses, systematically compare performance across tasks and model families, and identify error patterns or interesting behaviors. Table 4 illustrates the visualization of the example challenge, where we can immediately observe an unexpected behavior: adding web search tool on top of coding for GPT-5 (high) actually degrades the model performance for this particular challenge. The full interactive demo including all model responses is available at critpt.com/example.
@@ A.1 List of CritPt challenges
    Nieh-Yan modified gravity with Torsion in FRW space-time
  10. 10.
@@
    Scattering rate of the Hatsugai-Kohmoto model
@@
    Dark matter detection with Cosmic Explorer
@@
    LIGO with modified mirrors for ultralight vector dark matter
@@
    Magnetic space group identification
@@
    Charge density wave diffraction
@@
    Superresolution
@@
    Quantum games in multi-slit interference
@@
    Noise robustness in Kochen–Specker and Spekkens contextuality
@@
    Conformal correlators
@@
    Constructing fermionic matrix operators
@@
    Quantum capacity for quantum channels
@@
    Contraction coefficients for quantum relative entropy
@@
    Row-Twirling channel on qubit lattice
@@ A.4 API usage statistics
GPT-5 (high, code & web) 330569 1.25 291050 0.125 10600 10.0 2286 10.0 343455 0.578
GPT-5 (high, code) 17728 1.25 16539 0.125 11379 10.0 1996 10.0 31102 0.158
GPT-5 (high) 864 1.25 55 0.125 24686 10.0 2035 10.0 27585 0.268
o3 (high) 864 2.0 200 0.5 17356 8.0 1322 8.0 19653 0.151
o4-mini (high) 864 1.1 140 0.275 16096 4.4 1536 4.4 18496 0.079
Gemini 2.5 Pro 1012 1.25 0 0.31 20943 10.0 3196 10.0 25151 0.243
Gemini 2.5 Flash 1012 0.3 0 0.075 22031 2.5 3726 2.5 26769 0.065
DeepSeek R1 821 0.56 320 0.07 19656 1.68 1236 1.68 21713 0.036
Claude Opus 4 1005 15.0 0 1.5 4566 75.0 8637 75.0 14208 1.005
GPT-5 (minimal) 864 1.25 24 0.125 0 10.0 2315 10.0 3179 0.024
GPT-4o 865 2.5 216 1.25 0 10.0 950 10.0 1815 0.012
Llama-4 Maverick 818 0.27 0 0.0 0 0.85 1383 0.85 2201 0.001